OpenStudy (anonymous):

please, i really need help with this question. no one seems to know how to do it Find the values of x and y that maximize or minimize the objective function for each of these graphs. Minimum for P = 3x + y

OpenStudy (anonymous):

Umhh.. my Best Suggestion would be finiding a calculator :))

OpenStudy (anonymous):

But i have absolutely no idea where to start. can you help? or do you know who can?

OpenStudy (anonymous):

@lovablenene

OpenStudy (anonymous):

where is the graph ?

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

here it is @sara12345

OpenStudy (anonymous):

Okay, minimum/maximum value occurs at the corners of shaded region

OpenStudy (anonymous):

look at the graph, points A, B and C are the corner points of shaded region, right ?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

you need to evaluate the given objective function at these 3 corner points

OpenStudy (anonymous):

ok so what do i do to do that?

OpenStudy (anonymous):

P = 3x + y find the value of P , at A = (2, 1)

OpenStudy (anonymous):

put x = 2, y = 1 in P P = 3x + y = 3(2) + 1 = 7

OpenStudy (anonymous):

similarly, evaluate P at other two corner points B and C also

OpenStudy (anonymous):

3(2) +1 p=7?

OpenStudy (anonymous):

Yes 7 is its value at A, evaluate P at B = (3, 0)

OpenStudy (anonymous):

p=4

OpenStudy (anonymous):

no i mean 1

OpenStudy (anonymous):

P = 3x + y B = (3, 0) put x = 3, y = 0 in P, wat do u get ?

OpenStudy (anonymous):

oh ok x is 3(3)+0 p=9

OpenStudy (anonymous):

good, similarly evaluate P at C also, wat do u get?

OpenStudy (anonymous):

c is 0,5 so p=3(0)+5 p=5

OpenStudy (anonymous):

Yes ! so, you got, P value at A = 7 P value at B = 9 P value at C = 5

OpenStudy (anonymous):

ok that part was easy now what.

OpenStudy (anonymous):

We're done.

OpenStudy (anonymous):

Since, Max/Min value occurs at corner points only, the minimum of above values is the minimum value of objective function.

OpenStudy (anonymous):

really??? thank you so much!!!!! you really helped me through it! made a lot of sense ans was easy!!! thank you so,so much!!!!!

OpenStudy (anonymous):

that is, Minimum value for objective funciton occurs at C; and the minimum value is 5

OpenStudy (anonymous):

glad to hear :) yw !!

OpenStudy (anonymous):

Yayyy :))) I just Learned Something New Thanks Too Everyone Who Helped Me Out Too I Dont Understand Most Of This :))

OpenStudy (anonymous):